OpenStudy (anonymous):

F(x)= 3(x-3)(x+2)^2 A) using limits describe the "end behavior" B) State each zero, it's multiplicity, and it whether or not the graph crosses through it.

OpenStudy (irishboy123):

for \(F(x)= 3(x-3)(x+2)^2\) A) i'd used \(F = O(x^3)\) B) they've already factored it for you so just count the roots. useful rule of thumb: odd multiplicity crosses the axis but even multiplicity just touches the axis.....